On January 1, 2016, Tremblay Transport purchased a $160,000 truck

Tremblay plans on driving the truck for 400,000 kilometers.

Expected residual value for the truck is $40,000.

On June 30, 2019, after having been driven 180 000 kilometers, the truck had an accident on the highway and Tremblay had to sell it for $42,000 cash.

REQUIRED :

Record the disposal of the truck on June 30, 2019, assuming the amortization expense for the truck to decembre 31, 2018 has already been recorded using the unit-of-production method and that the truck has been driven 20 000 km since then.
Answer #1

Answer:

Date Account Titles and Explanation Jun 30, 19 Depreciation Expense To Accumulated depreciation (Being Depreciation applied t

Explanation:

Depreciation Expense(Current year till date of disposal)=(Original Cost-Salvage Value)*Kms Driven/Total Life in Kms=(160,000-40,000)*20,000/400,000=6,000

Total accumulated depreciation till date=(Original Cost-Salvage Value)*Kms Driven/Total Life in Kms=(160,000-40,000)*180,000/400,000=54,000

Loss on sale=(Original Cost-accumulated depreciation)-Sales price=(160,000-54,000)-42,000=64,000

If helpful, Thumbs UP please:)